# Vision Mamba: Efficient Visual Representation Learning with Bidirectional State Space Model

### 17 Jan 2024 v1

@@ Abstract
Recently the state space models (SSMs) with efficient hardware-aware designs, i.e., Mamba, have shown great potential for long sequence modeling. Building efficient and generic vision backbones purely upon SSMs is an appealing direction. However, representing visual data is challenging for SSMs due to the position-sensitivity of visual data and the requirement of global context for visual understanding. In this paper, we show that the reliance of visual representation learning on self-attention is not necessary and propose a new generic vision backbone with bidirectional Mamba blocks (Vim), which marks the image sequences with position embeddings and compresses the visual representation with bidirectional state space models. On ImageNet classification, COCO object detection, and ADE20k semantic segmentation tasks, Vim achieves higher performance compared to well-established vision transformers like DeiT, while also demonstrating significantly improved computation & memory efficiency. For example, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images with a resolution of 1248$\times$1248. The results demonstrate that Vim is capable of overcoming the computation & memory constraints on performing Transformer-style understanding for high-resolution images and it has great potential to become the next-generation backbone for vision foundation models.
Figure 1: Performance and efficiency comparisons between DeiT [60] and our Vim model. For the accuracy comparison, we first pretrain DeiT and Vim on IN1K classification dataset [10], then we finetune the generic backbones on different downstream dense prediction tasks, i.e., semantic segmentation, object detection, instance segmentation. Results show that the proposed Vim outperforms DeiT on both pretraining and finetuning tasks. For the speed comparison, the proposed Vim is significantly faster than DeiT when dealing with large images because of its subquadratic-time computation. For the GPU memory comparison, Vim requires less GPU memory than DeiT for inferring large images by its linear memory complexity. With not only superior accuracy for vision tasks, the proposed Vim is also more efficient than DeiT in dealing with large images. For example, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images with a resolution of 1248$\times$1248, i.e., 6084 tokens per image.
@@ 1 Introduction
Recent research advancements have led to a surge of interest in the state space model (SSM). Originating from the classic state space model [30], modern SSMs excel at capturing long-range dependencies and benefit from parallel training. Some SSM-based methods, such as the linear state-space layers (LSSL) [22], structured state space sequence model (S4) [21], diagonal state space (DSS) [24], and S4D [23], are proposed to process sequence data across a wide range of tasks and modalities, particularly on modeling long-range dependencies. They are efficient on long sequences because of convolutional computation and near-linear computation. 2-D SSM [2], SGConvNeXt [37], and ConvSSM [52] combine SSM with CNN or Transformer architecture to process 2-D data. The recent work, Mamba [20], incorporates time-varying parameters into the SSM and proposes a hardware-aware algorithm to enable efficient training and inference. The superior scaling performance of Mamba indicates that it is a promising alternative to Transformer in language modeling. Nevertheless, a generic pure-SSM-based backbone has not been explored for vision tasks.
@@
Vision Transformers (ViTs) have achieved great success in visual representation learning, excelling in both large-scale self-supervised pre-training and high performance on downstream tasks. Compared with convolutional neural networks, the core advantage lies in that ViT can provide each image patch with data/patch-dependent global context through self-attention. This differs from convolutional networks that use the same parameters, i.e., the convolutional filters, for all positions. Another advantage is the modality-agnostic modeling by treating an image as a sequence of patches without 2D inductive bias, which makes it the preferred architecture for multimodal applications [3, 36, 40]. At the same time, the self-attention mechanism in Transformers poses challenges in terms of speed and memory usage when dealing with long-range visual dependencies, e.g., processing high-resolution images.
@@
Motivated by the success of Mamba in language modeling, it is appealing that we can also transfer this success from language to vision, i.e., to design a generic and efficient visual backbone with the advanced SSM method. However, there are two challenges for Mamba, i.e., unidirectional modeling and lack of positional awareness. To address these challenges, we propose the Vision Mamba (Vim) block, which incorporates the bidirectional SSMs for data-dependent global visual context modeling and position embeddings for location-aware visual recognition. We first split the input image into patches and linearly project them as vectors to Vim. Image patches are treated as the sequence data in Vim blocks, which efficiently compresses the visual representation with the proposed bidirectional selective state space. Furthermore, the position embedding in Vim block provides the awareness for spatial information, which enables Vim to be more robust in dense prediction tasks. In the current stage, we train the Vim model on the supervised image classification task using the ImageNet dataset and then use the pretrained Vim as the backbone to perform sequential visual representation learning for downstream dense prediction tasks, i.e., semantic segmentation, object detection, and instance segmentation. Like Transformers, Vim can be pretrained on large-scale unsupervised visual data for better visual representation. Thanks to the better efficiency of Mamba, the large-scale pretraining of Vim can be achieved with lower computational cost.
@@
  • Without the need of attention, the proposed Vim has the same modeling power as ViT while it only has subquadratic-time computation and linear memory complexity. Specifically, our Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images at the resolution of 1248$\times$1248.
@@
  • Benefiting from the efficient hardware-aware design of Mamba, Vim is much more efficient than the self-attention-based DeiT [60] for high-resolution computer vision tasks, e.g., video segmentation, aerial image analysis, medical image segmentation, computational pathology.
@@ 2 Related Work
Architectures for generic vision backbone.
In the early eras, ConvNet [34] serves as the de-facto standard network design for computer vision. Many convolutional neural architectures [33, 56, 51, 25, 57, 63, 26, 72, 58, 50] have been proposed as the vision backbone for various visual applications. The pioneering work, Vision Transformer (ViT) [14] changes the landscape. It treats an image as a sequence of flattened 2D patches and directly applies a pure Transformer architecture. The surprising results of ViT on image classification and its scaling ability encourage a lot of follow-up works [61, 59, 62, 16]. One line of works focuses on hybrid architecture designs by introducing 2D convolutional priors into ViT [69, 9, 15, 13]. PVT [66] proposes a pyramid structure Transformer. Swin Transformer [42] applies self-attention within shift windows. Another line of works focuses on improving traditional 2D ConvNets with more advanced settings [67, 41]. ConvNeXt [43] reviews the design space and proposes pure ConvNets, which can be scalable as ViT and its variants. RepLKNet [12] proposes to scale up the kernel size of existing ConvNets to bring improvements.
@@
State space models for long sequence modeling.
[21] proposes a Structured State-Space Sequence (S4) model, a novel alternative to CNNs or Transformers, to model the long-range dependency. The promising property of linearly scaling in sequence length attracts further explorations. [53] proposes a new S5 layer by introducing MIMO SSM and efficient parallel scan into S4 layer. [18] designs a new SSM layer, H3, that nearly fills the performance gap between SSMs and Transformer attention in language modeling. [46] builds the Gated State Space layer on S4 by introducing more gating units to improve the expressivity. Recently, [20] proposes a data-dependent SSM layer and builds a generic language model backbone, Mamba, which outperforms Transformers at various sizes on large-scale real data and enjoys linear scaling in sequence length. In this work, we explore transferring the success of Mamba to vision, i.e., building a generic vision backbone purely upon SSM without attention.
State space models for visual applications.
[27] uses 1D S4 to handle the long-range temporal dependencies for video classification. [47] further extends 1D S4 to handle multi-dimensional data including 2D images and 3D videos. [28] combines the strengths of S4 and self-attention to build TranS4mer model, achieving state-of-the-art performance for movie scene detection. [64] introduces a novel selectivity mechanism to S4, largely improving the performance of S4 on long-form video understanding with a much lower memory footprint. [73] supplants attention mechanisms with a more scalable SSM-based backbone to generate high-resolution images and process fine-grained representation under affordable computation. [45] proposes U-Mamba, a hybrid CNN-SSM architecture, to handle the long-range dependencies in biomedical image segmentation. The above works either apply SSM to specific visual applications or build a hybrid architecture by combining SSM with convolution or attention. Different from them, we build a pure-SSM-based model, which can be adopted as a generic vision backbone.
@@ 3.3 Vim Block
Input: token sequence $\mathbf{T}_{l-1}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$
Output: token sequence $\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$
/* normalize the input sequence $\mathbf{T}_{l-1}^{\prime}$/
$\mathbf{T}_{l-1}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$leftarrow $\mathbf{Norm}(\mathbf{T}_{l-1})$ $\mathbf{x}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{x}}(\mathbf{T}_{l-1}^{\prime})$ $\mathbf{z}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{z}}(\mathbf{T}_{l-1}^{\prime})$ /* process with different direction */
1 for $o$ in {forward, backward} do
       $\mathbf{x}^{\prime}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{SiLU}(\mathbf{Conv1d}_{o}(\mathbf{x}))$ $\mathbf{B}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{B}}_{o}(\mathbf{x}^{\prime}_{o})$ $\mathbf{C}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{C}}_{o}(\mathbf{x}^{\prime}_{o})$ /* softplus ensures positive $\mathbf{\Delta}_{o}$/
       $\mathbf{\Delta}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\log(1+\exp(\mathbf{Linear}^{\mathbf{\Delta}}_{o}(\mathbf{x}^{\prime}_{o})+\mathbf{Parameter}^{\mathbf{\Delta}}_{o}))$ /* shape of $\mathbf{Parameter}^{\mathbf{A}}_{o}$ is $(\mathtt{E},\mathtt{N})$/
2       $\overline{\mathbf{A}}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{Parameter}^{\mathbf{A}}_{o}$ $\overline{\mathbf{B}}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{B}_{o}$ $\mathbf{y}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{SSM}(\overline{\mathbf{A}}_{o},\overline{\mathbf{B}}_{o},\mathbf{C}_{o})(\mathbf{x}_{o}^{\prime})$
3 end for
/* get gated $\mathbf{y}_{o}$/
$\mathbf{y}_{forward}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{forward}\bigodot\mathbf{SiLU}(\mathbf{z})$ $\mathbf{y}_{backward}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{backward}\bigodot\mathbf{SiLU}(\mathbf{z})$ /* residual connection */
$\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$leftarrow $\mathbf{Linear}^{\mathbf{T}}(\mathbf{y}_{forward}^{\prime}+\mathbf{y}_{backward}^{\prime})+\mathbf{T}_{l-1}$ Return: $\mathbf{T}_{l}$
@@
Specifically, we present the operations of Vim block in Algo. 1. The input token sequence $\mathbf{T}_{\mathtt{l}-1}$ is first normalized by the normalization layer. Next, we linearly project the normalized sequence to the $\mathbf{x}$ and $\mathbf{z}$ with dimension size $E$. Then, we process the $\mathbf{x}$ from the forward and backward directions. For each direction, we first apply the 1-D convolution to the $\mathbf{x}$ and get the $\mathbf{x}^{\prime}_{o}$. We then linearly project the $\mathbf{x}^{\prime}_{o}$ to the $\mathbf{B}_{o}$, $\mathbf{C}_{o}$, $\mathbf{\Delta}_{o}$, respectively. The $\mathbf{\Delta}_{o}$ is then used to transform the $\overline{\mathbf{A}}_{o}$, $\overline{\mathbf{B}}_{o}$, respectively. Finally, we compute the $\mathbf{y}_{forward}$ and $\mathbf{y}_{backward}$ through the SSM. The $\mathbf{y}_{forward}$ and $\mathbf{y}_{backward}$ are then gated by the $\mathbf{z}$ and added together to get the output token sequence $\mathbf{T}_{\mathtt{l}}$.

### 10 Feb 2024 v2

@@ Abstract
Recently the state space models (SSMs) with efficient hardware-aware designs, i.e., the Mamba deep learning model, have shown great potential for long sequence modeling. Meanwhile building efficient and generic vision backbones purely upon SSMs is an appealing direction. However, representing visual data is challenging for SSMs due to the position-sensitivity of visual data and the requirement of global context for visual understanding. In this paper, we show that the reliance on self-attention for visual representation learning is not necessary and propose a new generic vision backbone with bidirectional Mamba blocks (Vim), which marks the image sequences with position embeddings and compresses the visual representation with bidirectional state space models. On ImageNet classification, COCO object detection, and ADE20k semantic segmentation tasks, Vim achieves higher performance compared to well-established vision transformers like DeiT, while also demonstrating significantly improved computation & memory efficiency. For example, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images with a resolution of 1248$\times$1248. The results demonstrate that Vim is capable of overcoming the computation & memory constraints on performing Transformer-style understanding for high-resolution images and it has great potential to be the next-generation backbone for vision foundation models.
Figure 1: Performance and efficiency comparisons between DeiT [59] and our Vim model. For the accuracy comparison, we first pretrain DeiT and Vim on IN1K classification dataset [9], then we finetune the generic backbones on different downstream dense prediction tasks, i.e., semantic segmentation, object detection, instance segmentation. Results show that the proposed Vim outperforms DeiT on both pretraining and finetuning tasks. Vim is also more computation and memory efficient than DeiT in dealing with high-resolution images. For example, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images with a resolution of 1248$\times$1248, i.e., 6084 tokens per image.
@@ 1 Introduction
Recent research advancements have led to a surge of interest in the state space model (SSM). Originating from the classic Kalman filter model [29], modern SSMs excel at capturing long-range dependencies and benefit from parallel training. Some SSM-based methods, such as the linear state-space layers (LSSL) [21], structured state space sequence model (S4) [20], diagonal state space (DSS) [23], and S4D [22], are proposed to process sequence data across a wide range of tasks and modalities, particularly on modeling long-range dependencies. They are efficient in processing long sequences because of convolutional computation and near-linear computation. 2-D SSM [2], SGConvNeXt [36], and ConvSSM [51] combine SSM with CNN or Transformer architecture to process 2-D data. The recent work, Mamba [19], incorporates time-varying parameters into the SSM and proposes a hardware-aware algorithm to enable very efficient training and inference. The superior scaling performance of Mamba indicates that it is a promising alternative to Transformer in language modeling. Nevertheless, a generic pure-SSM-based backbone network has not been explored for processing visual data, such as images and videos.
@@
Vision Transformers (ViTs) have achieved great success in visual representation learning, excelling in large-scale self-supervised pre-training and high performance on downstream tasks. Compared with convolutional neural networks, the core advantage lies in that ViT can provide each image patch with data/patch-dependent global context through self-attention. This differs from convolutional networks that use the same parameters, i.e., the convolutional filters, for all positions. Another advantage is the modality-agnostic modeling by treating an image as a sequence of patches without 2D inductive bias, which makes it the preferred architecture for multimodal applications [3, 35, 39]. At the same time, the self-attention mechanism in Transformers poses challenges in terms of speed and memory usage when dealing with long-range visual dependencies, e.g., processing high-resolution images.
@@
Motivated by the success of Mamba in language modeling, it is appealing that we can also transfer this success from language to vision, i.e., to design a generic and efficient visual backbone with the advanced SSM method. However, there are two challenges for Mamba, i.e., unidirectional modeling and lack of positional awareness. To address these challenges, we propose the Vision Mamba (Vim) model, which incorporates the bidirectional SSMs for data-dependent global visual context modeling and position embeddings for location-aware visual recognition. We first split the input image into patches and linearly project them as vectors to Vim. Image patches are treated as the sequence data in Vim blocks, which efficiently compresses the visual representation with the proposed bidirectional selective state space. Furthermore, the position embedding in Vim block provides the awareness for spatial information, which enables Vim to be more robust in dense prediction tasks. In the current stage, we train the Vim model on the supervised image classification task using the ImageNet dataset and then use the pretrained Vim as the backbone to perform sequential visual representation learning for downstream dense prediction tasks, i.e., semantic segmentation, object detection, and instance segmentation. Like Transformers, Vim can be pretrained on large-scale unsupervised visual data for better visual representation. Thanks to the better efficiency of Mamba, the large-scale pretraining of Vim can be achieved with lower computational cost.
@@
  • Without the need of attention, the proposed Vim has the same modeling power as ViT while it only has subquadratic-time computation and linear memory complexity. Specifically, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images at the resolution of 1248$\times$1248.
@@ 2 Related Work
Architectures for generic vision backbone. In the early eras, ConvNet [33] serves as the de-facto standard network design for computer vision. Many convolutional neural architectures [32, 55, 50, 24, 56, 62, 25, 71, 57, 49] have been proposed as the vision backbone for various visual applications. The pioneering work, Vision Transformer (ViT) [13] changes the landscape. It treats an image as a sequence of flattened 2D patches and directly applies a pure Transformer architecture. The surprising results of ViT on image classification and its scaling ability encourage a lot of follow-up works [60, 58, 61, 15]. One line of works focuses on hybrid architecture designs by introducing 2D convolutional priors into ViT [68, 8, 14, 12]. PVT [65] proposes a pyramid structure Transformer. Swin Transformer [41] applies self-attention within shift windows. Another line of works focuses on improving traditional 2D ConvNets with more advanced settings [66, 40]. ConvNeXt [42] reviews the design space and proposes pure ConvNets, which can be scalable as ViT and its variants. RepLKNet [11] proposes to scale up the kernel size of existing ConvNets to bring improvements.
@@
State space models for long sequence modeling. [20] proposes a Structured State-Space Sequence (S4) model, a novel alternative to CNNs or Transformers, to model the long-range dependency. The promising property of linearly scaling in sequence length attracts further explorations. [52] proposes a new S5 layer by introducing MIMO SSM and efficient parallel scan into S4 layer. [17] designs a new SSM layer, H3, that nearly fills the performance gap between SSMs and Transformer attention in language modeling. [45] builds the Gated State Space layer on S4 by introducing more gating units to improve the expressivity. Recently, [19] proposes a data-dependent SSM layer and builds a generic language model backbone, Mamba, which outperforms Transformers at various sizes on large-scale real data and enjoys linear scaling in sequence length. In this work, we explore transferring the success of Mamba to vision, i.e., building a generic vision backbone purely upon SSM without attention.
State space models for visual applications. [26] uses 1D S4 to handle the long-range temporal dependencies for video classification. [46] further extends 1D S4 to handle multi-dimensional data including 2D images and 3D videos. [27] combines the strengths of S4 and self-attention to build TranS4mer model, achieving state-of-the-art performance for movie scene detection. [63] introduces a novel selectivity mechanism to S4, largely improving the performance of S4 on long-form video understanding with a much lower memory footprint. [72] supplants attention mechanisms with a more scalable SSM-based backbone to generate high-resolution images and process fine-grained representation under affordable computation. [44] proposes U-Mamba, a hybrid CNN-SSM architecture, to handle the long-range dependencies in biomedical image segmentation. The above works either apply SSM to specific visual applications or build a hybrid architecture by combining SSM with convolution or attention. Different from them, we build a pure-SSM-based model, which can be adopted as a generic vision backbone.
@@ 3.3 Vim Block
0:  token sequence $\mathbf{T}_{l-1}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$
0:  token sequence $\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$
1:  /* normalize the input sequence $\mathbf{T}_{l-1}^{\prime}$/
2:  $\mathbf{T}_{l-1}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$leftarrow $\mathbf{Norm}(\mathbf{T}_{l-1})$
3:  $\mathbf{x}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{x}}(\mathbf{T}_{l-1}^{\prime})$
4:  $\mathbf{z}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{z}}(\mathbf{T}_{l-1}^{\prime})$
5:  /* process with different direction */
6:  for $o$ in {forward, backward} do
7:     $\mathbf{x}^{\prime}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{SiLU}(\mathbf{Conv1d}_{o}(\mathbf{x}))$
8:     $\mathbf{B}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{B}}_{o}(\mathbf{x}^{\prime}_{o})$
9:     $\mathbf{C}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{C}}_{o}(\mathbf{x}^{\prime}_{o})$
10:     /* softplus ensures positive $\mathbf{\Delta}_{o}$/
11:     $\mathbf{\Delta}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\log(1+\exp(\mathbf{Linear}^{\mathbf{\Delta}}_{o}(\mathbf{x}^{\prime}_{o})+\mathbf{Parameter}^{\mathbf{\Delta}}_{o}))$
12:     /* shape of $\mathbf{Parameter}^{\mathbf{A}}_{o}$ is $(\mathtt{E},\mathtt{N})$/
13:     $\overline{\mathbf{A}}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{Parameter}^{\mathbf{A}}_{o}$
14:     $\overline{\mathbf{B}}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{B}_{o}$
15:     $\mathbf{y}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{SSM}(\overline{\mathbf{A}}_{o},\overline{\mathbf{B}}_{o},\mathbf{C}_{o})(\mathbf{x}_{o}^{\prime})$
16:  end for
17:  /* get gated $\mathbf{y}_{o}$/
18:  $\mathbf{y}_{forward}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{forward}\bigodot\mathbf{SiLU}(\mathbf{z})$
19:  $\mathbf{y}_{backward}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{backward}\bigodot\mathbf{SiLU}(\mathbf{z})$
20:  /* residual connection */
21:  $\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$leftarrow $\mathbf{Linear}^{\mathbf{T}}(\mathbf{y}_{forward}^{\prime}+\mathbf{y}_{backward}^{\prime})+\mathbf{T}_{l-1}$ Return: $\mathbf{T}_{l}$
Specifically, we present the operations of Vim block in Algo. 21. The input token sequence $\mathbf{T}_{\mathtt{l}-1}$ is first normalized by the normalization layer. Next, we linearly project the normalized sequence to the $\mathbf{x}$ and $\mathbf{z}$ with dimension size $E$. Then, we process the $\mathbf{x}$ from the forward and backward directions. For each direction, we first apply the 1-D convolution to the $\mathbf{x}$ and get the $\mathbf{x}^{\prime}_{o}$. We then linearly project the $\mathbf{x}^{\prime}_{o}$ to the $\mathbf{B}_{o}$, $\mathbf{C}_{o}$, $\mathbf{\Delta}_{o}$, respectively. The $\mathbf{\Delta}_{o}$ is then used to transform the $\overline{\mathbf{A}}_{o}$, $\overline{\mathbf{B}}_{o}$, respectively. Finally, we compute the $\mathbf{y}_{forward}$ and $\mathbf{y}_{backward}$ through the SSM. The $\mathbf{y}_{forward}$ and $\mathbf{y}_{backward}$ are then gated by the $\mathbf{z}$ and added together to get the output token sequence $\mathbf{T}_{\mathtt{l}}$.